Using stoichiometry, determine the mass of powdered drink mix needed to make a 1.0 M solution of 100 mL.
formula:(C12H22O11)
1
Expert's answer
2019-05-28T06:21:24-0400

Molar concentration can be determined by the following equation:

"c=\frac {n}{V}"

, where n - amount in moles, V - volume of solution in liters.

"n=\frac{m}{M}"

, where m - mass of powdered drink mix, and M - molar mass of powdered drink mix.

Thus, "c=\frac{m}{M \cdot V}"

"M(C_{12}H_{22}O_{11})=12 \cdot 12 \frac{g}{mol} + 22 \cdot 1 \frac{g}{mol} + 11 \cdot 16 \frac{g}{mol} = 342 \frac{g}{mol}"

"m = M \cdot V \cdot c = 342 \frac {g}{mol} \cdot 0.1 L \cdot 1.0 \frac{mol}{L} = 34.2 g"